### v1

# Coupling Brownian loop soups and random walk loop soups at all polynomial scales

Abstract
Lawler and Trujillo Ferreras [4] constructed a well-known coupling between the Brownian loop soups in $\mathbb{R}^{2}$ and the random walk loop soups on $\mathbb{Z}^{2}$ (one rescales the random walk loops by $1/N$, their time parametrizations by $1/(2N^{2})$, and let $N\to\infty$), which led to numerous applications. It nevertheless only holds for loops with time length at least $N^{\theta-2}$ for $\theta\in(2/3,2)$. In particular, there is no control on mesoscopic loops with time length less than $N^{-4/3}$ (i.e. roughly diameter less than $N^{-2/3}$).
In this paper, we find a simple way to remove the restriction $\theta>2/3$, so that such a coupling works for all $\theta\in(0,2)$, i.e. for loops at all polynomial scales. We also establish this coupling in any dimension $d\geq 1$ (i.e. for random walk loop soups on $\mathbb{Z}^{d}$ and Brownian loop soups on $\mathbb{R}^{d}$).
1 Introduction
The Brownian loop soup, introduced by Lawler and Werner [5], plays a prominent role in random geometry. The random walk loop soup, introduced by Lawler and Trujillo Ferreras [4], is the discrete analogue of the Brownian loop soup. It was proved in [4] that the random walk loop soup on the square lattice $\mathbb{Z}^{2}$ converges to the Brownian loop soup in $\mathbb{R}^{2}$ in the scaling limit. Moreover, a precise coupling between the two loop soups was constructed, as we state below.
1.1 Lawler and Trujillo Ferreras’s coupling
Let $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ be respectively a realization of the Brownian loop soup on $\mathbb{R}^{2}$ and random walk loop soup on $\mathbb{Z}^{2}$ with intensity $\lambda$ (see Section 2 for precise definitions). Throughout this paper, we regard both the Brownian loops and random walk loops as rooted loops. We also consider the random walk loops as curves by linear interpolation.
For each positive integer $N$, we define $\mathcal{A}_{\lambda,N}$ to be the collection of loops obtained from $\mathcal{A}_{\lambda}$ by scaling space by $1/N$ and time by $1/N^{2}$. More precisely, for each loop $\gamma$ with time length $t_{\gamma}$, let $\Phi_{N}\gamma$ be the loop with
$t_{\Phi_{N}\gamma}=t_{\gamma}/N^{2},\qquad\Phi_{N}\gamma(t)=N^{-1}\gamma(tN^{2}),\quad 0\leq t\leq t_{\gamma}/N^{2}.$ (1.1)
We then set
$\mathcal{A}_{\lambda,N}=\{\Phi_{N}\gamma:\gamma\in\mathcal{A}_{\lambda}\}.$
The scaling invariance of the Brownian loop soup implies that $\mathcal{A}_{\lambda,N}$ is also distributed as a Brownian loop soup with intensity $\lambda$.
For the random walk loop soup, we define $\widetilde{\Phi}_{N}\gamma$ to be the loop with
$t_{\widetilde{\Phi}_{N}\gamma}=t_{\gamma}/(2N^{2}),\qquad\widetilde{\Phi}_{N}\gamma(t)=N^{-1}\gamma(t2N^{2}),\quad 0\leq t\leq t_{\gamma}/(2N^{2}),$ (1.2)
and let
$\widetilde{\mathcal{A}}_{\lambda,N}=\{\widetilde{\Phi}_{N}\gamma:\gamma\in\widetilde{\mathcal{A}}_{\lambda}\}.$ (1.3)
As pointed out in [4], the extra factor $2$ in the time scaling for the random walk loops comes from the fact that the covariance of simple two-dimensional random walk in $2n$ steps is $nI$ as opposed to $2nI$ for a Brownian motion at time $2n$.
To state the coupling by Lawler and Trujillo Ferreras, we also need to define the following auxiliary functions. For $t\geq(5/8)N^{-2}$ and integer $k\geq 1$, let
$\varphi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{k-3/8}{N^{2}}\leq t<\frac{k+5/8}{N^{2}}.$
For $z\in\mathbb{C}$ and $z_{0}\in\mathbb{Z}^{2}$, define
$\psi_{N}(z)=\frac{z_{0}}{N}\quad\text{if}\quad\max\{|\Re\{Nz-z_{0}\}|,|\Im\{Nz-z_{0}\}|\}<1/2.$ (1.4)
Then definition of $\psi_{N}(z)$ when $Nz$ is exactly on an edge of the dual lattice of $\mathbb{Z}^{2}$ is irrelevant for the following theorem.
Theorem 1.1 (Theorem 1.1, [4]).
One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$, $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{2}$, and $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the random walk loop soup on $\mathbb{Z}^{2}$. Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
Moreover, for $\mathcal{A}_{\lambda,N}$ and $\widetilde{\mathcal{A}}_{\lambda,N}$ defined as above, there exists $c>0$ such that for every $r\geq 1$, $\lambda>0$, integer $N\geq 1$ and every $\theta\in(2/3,2)$, except perhaps on an event of probability at most $c(\lambda+1)r^{2}N^{2-3\theta}$, there is a one-to-one correspondence between the following two sets of loops
$\{\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}:t_{\widetilde{\gamma}}>N^{\theta-2},|\widetilde{\gamma}(0)|<r\},\quad\{\gamma\in\mathcal{A}_{\lambda,N}:\varphi_{N}(t_{\gamma})>N^{\theta-2},|\psi_{N}(\gamma(0))|<r\}.$
If $\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}$ and $\gamma\in\mathcal{A}_{\lambda,N}$ are paired in this correspondence, then
$|t_{\gamma}-t_{\widetilde{\gamma}}|\leq 5/8N^{-2},\qquad$
$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$
The coupling by Lawler and Trujillo Ferreras has proved to be useful in a great number of situations, allowing one to transfer the information about the Brownian loop soup back to the random walk loop soup, and vice versa. However, since we can only take $\theta>2/3$ in Theorem 1.1, this coupling does not control mesoscopic loops with time length less than $N^{-4/3}$ (which roughly corresponds to diameter less than $N^{-2/3}$). In practice, it is often necessary to control the mesoscopic loops, since they can potentially play a role in the scaling limit in terms of connectivity. For instance, the restriction $\theta>2/3$ was problematic for the purpose of Aïdékon, Berestycki, Jego and Lupu [1], when approximating the multiplicative chaos of the Brownian loop soup by that of the random walk loop soup (see Sections 1.5 and 11.3 in [1] for more details). In [1, Lemma 11.11], they constructed a coupling for mesoscopic loops at all polynomial scales, with the drawback that it only works for local loops, i.e. the loops whose roots are close to a given point $z$ (roughly, the distance of the root of a loop to $z$ is at most the order of the diameter of that loop).
1.2 Main result
It is unclear at first sight whether $\theta>2/3$ can be removed, [^1] For example, this was believed to be imposssible in [1, Section 11.3]: “On the other hand, it is fairly clear from the proof of [4] that their result is sharp, and that the coupling described above cannot hold without the restriction $\theta>2/3$; that is, at all scales smaller than $N^{-2/3}$ some discrete and continuous loops somewhere will be quite different from one another.” since the value $2/3$ comes from a computation involving the following Taylor expansion of the total mass $q_{2}(n)$ of all random walk loops on $\mathbb{Z}^{2}$ with length $2n$ rooted at 0
$\widetilde{q}_{2}(n)=(2n)^{-1}\left[2^{-2n}\binom{2n}{n}\right]^{2}=\frac{1}{2\pi n^{2}}-\frac{1}{8\pi n^{3}}+O(n^{-4}).$ (1.5)
Somewhat unexpectedly, after a close study of the proof by Lawler and Trujillo Ferreras [4], we have found a natural way to get rid of the constraint $\theta>2/3$. This shows that the value $2/3$ in fact does not play an inherent role, and the size $N^{-2/3}$ is not an impassable barrier. Our proof follows the outline of [4], by carefully tuning some parameters. Our observation is rather simple, but it seems that it has been overlooked.
We will directly establish our result for random walks on $\mathbb{Z}^{d}$ for all $d\geq 1$. Note that the combinatorial formula (1.5) relies on a trick which is specific for $d=2$, namely we can rotate $\mathbb{Z}^{2}$ by $45$ degrees, so that the original simple random walk projected to the vertical and horizontal axes are independent simple random walks. For $d\geq 3$, it is more convenient to consider the continuous-time simple random walk, which spends an (independent) exponential time with mean $1$ at each vertex. We will rely on the following Taylor development of the total mas $q_{d}(t)$ of all continuous-time random walk loops on $\mathbb{Z}^{d}$ with time length $t$ rooted at 0 (see (2.3) and (2.4))
$q_{d}(t)=(d/2)^{d/2}\pi^{-d/2}t^{-d/2-1}+O\big(t^{-d/2-2}\big).$ (1.6)
Our proof requires only the first term in the Taylor series, unlike (1.5).
In order to state our result, we keep the same notations as in Section 1.1, with the following obvious extensions. For $d\geq 1$, for a simple random walk loop $\gamma$ on $\mathbb{Z}^{d}$, let $\widetilde{\Phi}_{N}\gamma$ be the loop with
$t_{\widetilde{\Phi}_{N}\gamma}=t_{\gamma}/(dN^{2}),\qquad\widetilde{\Phi}_{N}\gamma(t)=N^{-1}\gamma(tdN^{2}),\quad 0\leq t\leq t_{\gamma}/(dN^{2}).$
For $z\in\mathbb{R}^{d}$ and $z_{0}\in\mathbb{Z}^{d}$, let $\psi_{N}(z)=z_{0}/N$, if $z$ is in the unit hypercube centered at $z_{0}$. The definition of $\psi_{N}(z)$ when $z$ falls on the boundary of any unit hypercube centered at a vertex of $\mathbb{Z}^{d}$ is irrelevant for our theorem.
The main new input in our proof is the introduction of a positive increasing sequence $\{a_{n}\}_{n\geq 1}$, whose precise definition is given in (3.2). We also show (in Lemma 3.1) that there exists a constant $c(d)>0$, such that
$|a_{n}-2n/d|\leq c(d).$ (1.7)
For $t\geq a_{1}N^{-2}$ and integer $k\geq 1$, let
$\chi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{a_{k}}{N^{2}}\leq t<\frac{a_{k+1}}{N^{2}}.$
Theorem 1.2.
Fix $d\geq 1$. One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$,
  1. (1)
    $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{d}$;
  2. (2)
    $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the continuous-time random walk loop soup on $\mathbb{Z}^{d}$ (or $\widetilde{\mathcal{A}}_{\lambda}$ can also be a discrete-time random walk loop soup if $d=1,2$).
Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$. Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
$\{\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}:t_{\widetilde{\gamma}}>N^{\theta-2},|\widetilde{\gamma}(0)|<r\},\quad\{\gamma\in\mathcal{A}_{\lambda,N}:\chi_{N}(t_{\gamma})>N^{\theta-2},|\psi_{N}(\gamma(0))|<r\}.$
If $\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}$ and $\gamma\in\mathcal{A}_{\lambda,N}$ are paired in this correspondence, then for the constant $c(d)$ in (1.7),
$|t_{\gamma}-t_{\widetilde{\gamma}}|\leq c(d)N^{-2},\qquad$ (1.8)
$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$ (1.9)
This strengthened coupling, which controls loops at any polynomial scale, rooted anywhere in a macroscopic domain, can have many potential applications. For example, in the two-dimensional case, it was proved by [6, 9] that the outer boundaries of the outermost clusters of the random walk loop soups with intensity $\lambda\in(0,1/2]$ converge to those of the Brownian loop soups (which are in turn distributed as conformal loop ensembles [8]). This convergence is nevertheless qualitative. If one wants to obtain a precise convergence rate, one may need to control the random walk loops at mesoscopic scales, everywhere in the domain simultaneously, since the outer boundary of a cluster is a global object.
On the other hand, for dimensions $d\geq 3$, to our knowledge, the convergence of random walk loop soups in $\mathbb{Z}^{d}$ to Brownian loop soups has not previously been written down, so we fill in this small gap in the literature. Note that it is clear that the same coupling also works for Brownian loop soups on the metric graph induced by $\mathbb{Z}^{d}$ (since its restriction to the vertices of $\mathbb{Z}^{d}$ is exactly a continuous-time random walk loop soup). This type of convergence has been used in e.g. [10, 7].
2 Preliminaries
2.1 Brownian loop soup
The Brownian loop soup in $\mathbb{R}^{d}$ with intensity $\lambda>0$ is a Poisson point process with intensity $\lambda\mu^{\rm br}$, where $\mu^{\rm br}$ is a $\sigma$-finite measure on Brownian loops defined as follows
$\mu^{\rm br}=\int_{\mathbb{R}^{d}}\int_{0}^{\infty}\frac{1}{t(2\pi t)^{d/2}}\mathbb{P}^{\rm br}(z,z;t)dtdz,$ (2.1)
where $dz$ is according to the Lebesgue measure on $\mathbb{R}^{d}$, and $\mathbb{P}^{\rm br}(z,z;t)$ is the probability measure of a $d$-dimensional Brownian bridge (which is a loop) starting and ending at $z$ with time length $t$. The point $z$ is also called the root of the loop.
It is common to consider the loops in the Brownian loop soup as unrooted loops, by forgetting its root. However, for the purpose of this paper, we will always regard the Brownian loops as rooted loops, which makes it easier to compare the difference between two parametrized loops.
It was shown in [5] that the two-dimensional Brownian loop soup (as a set of unrooted loops) is invariant under conformal maps (up to time reparametrization). In this paper, we will use the fact that the $d$-dimensional Brownian loop soup (as a set of rooted loops) is invariant under the usual Brownian scaling (i.e. we scale space by $c$ and time by $c^{2}$), which is obvious from (2.1).
2.2 Discrete-time random walk loop soup
A rooted discrete-time random walk loop on the lattice $\mathbb{Z}^{d}$ is a path $(z_{1},\cdots,z_{j})$ in $\mathbb{Z}^{d}$ such that $|z_{i}-z_{i+1}|=1$ for all $1\leq i\leq j-1$ and $z_{j}=z_{1}$. Let $t_{\gamma}:=j-1$ be the length of $\gamma$. The point $z_{1}=z_{j}$ is called the root of the loop. One can also unroot the loops on $\mathbb{Z}^{d}$, but we will only look at rooted loops in this paper.
Let $\mu^{\rm drw}$ be the random walk loop measure which assigns weight $(2n)^{-1}(2d)^{-2n}$ to each rooted loop on $\mathbb{Z}^{d}$ with length $2n$, for any integer $n\geq 1$. For $z\in\mathbb{Z}^{d}$ and integer $n\geq 1$, let $\nu_{n,z}$ be the measure $\mu^{\rm drw}$ restricted to loops rooted at $z$ with length $2n$. Then we have
$\mu^{\rm drw}=\sum_{z\in\mathbb{Z}^{d}}\sum_{n\geq 1}\nu_{n,z}.$
The (discrete-time) random walk loop soup with intensity $\lambda$ is a Poisson point process with intensity $\lambda\mu^{\rm drw}$, which is a multi-set of random walk loops. Let (note that $|\nu_{n,z}|$ is the same for all $z$)
$\widetilde{q}_{d}(n):=|\nu_{n,z}|.$
For $d=1,2$, we have the exact combinatorial formula (1.5) for $\widetilde{q}_{2}(n)$, and
$\widetilde{q}_{1}(n)=(2n)^{-1}2^{-2n}\binom{2n}{n}$. (2.2)
2.3 Continuous-time random walk loop soup
Consider the continuous-time random walk on $\mathbb{Z}^{d}$ which spends en exponential time with mean $1$ at each vertex, before jumping to one of the neighboring vertices with probability $1/(2d)$. For $x,y\in\mathbb{Z}^{d}$ and $t>0$, let $p_{t}(x,y)$ be its transition kernel. Let $\mathbb{P}^{\rm crw}(x,y;t)$ be the probability measure of a continuous-time random walk bridge with time length $t$ starting at $x$ and ending at $y$. Let $\mu^{\rm crw}$ be the measure on random walk loops defined by
$\mu^{\rm crw}=\sum_{z\in\mathbb{Z}^{d}}\int_{0}^{\infty}\frac{1}{t}p_{t}(x,y)\mathbb{P}^{\rm crw}(z,z;t)dt.$
The (continuous-time) random walk loop soup with intensity $\lambda$ is a Poisson point process with intensity $\lambda\mu^{\rm crw}$. Let
$q_{d}(t):=\frac{1}{t}p_{t}(0,0).$ (2.3)
Thanks to the following Taylor development of $p_{t}(0,0)$ (see e.g. [3, Theorem 2.13])
$p_{t}(0,0)=\bigg(\frac{d}{2\pi t}\bigg)^{d/2}(1+O(t^{-1})),$ (2.4)
we get the Taylor development (1.6) of $q_{d}(t)$. We also define the following quantity
$Q_{d}(n):=\int_{2n}^{2n+2}q_{d}(t)dt.$ (2.5)
2.4 KMT coupling
To couple a Brownian loop closely to a random walk loop, we rely on a variant of the classical coupling by Komlós, Major, and Tusnády [2]. Recall that the KMT coupling works for a wide class of random walks, and a comprehensive review of it can be found in [3, Chapter 7].
In particular, for any aperiodic irreducible walk $S_{n}=X_{1}+\cdots+X_{n}$ on $\mathbb{Z}$ such that there exist $\sigma,b>0$ with
$\mathbb{E}[X_{1}]=0,\quad\mathbb{E}[X_{1}^{2}]=\sigma^{2},\quad\mathbb{E}[\exp(b|X_{1}|)]<\infty,$ (2.6)
we can couple it with a Brownian motion (see e.g. [3, Theorem 7.1.1]).
Theorem 2.1.
Suppose that $(S_{n})_{n\geq 0}$ is an aperiodic irreducible random walk on $\mathbb{Z}$ that satisfies (2.6). Then one can define on the same probability space a Brownian motion $B_{t}$ with variance $\sigma^{2}$ and $(S_{n})_{n\geq 0}$ such that the following holds. For each $\alpha<\infty$, there is a $c_{\alpha}$ such that
$\mathbb{P}\left[\max_{1\leq j\leq n}|S_{j}-B_{j}|\geq c_{\alpha}\log n\right]\leq c_{\alpha}n^{-\alpha}$
The idea of the proof of Theorem 2.1 has been used in [4] to prove the following coupling between a discrete-time simple random walk bridge on $\mathbb{Z}$ a Brownian bridge with length $2n$ (see [4, Theorem 6.4]). Note that even though the random walk bridges in Theorem 2.2 and Lemma 2.3 below are defined on discrete time and space, we still consider them as continuous curves parametrized by continuous time intervals, by linear interpolation. This is, however, different from continuous-time random walks.
Theorem 2.2.
There exist $0<c,\alpha<\infty$, such that for every positive integer $n$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a simple random walk bridge $S$ on $\mathbb{Z}$, both with length $2n$, such that for all $r>0$,
$\mathbb{P}\left[\sup_{0\leq t\leq 2n}|B_{t}-S_{t}|>rc\log n\right]\leq cn^{\alpha-r}.$
By the trick of rotating $\mathbb{Z}^{2}$ by $45$ degrees (see the second paragraph of Section 1.2), the simple random walk on $\mathbb{Z}^{2}$ is the combination of two independent simple random walks on $\mathbb{Z}$, hence Theorem 2.2 implies the following coupling (see [4, Lemma 3.1] and the remark just below it).
Lemma 2.3.
There exists a constant $c_{0}>0$, such that for each $\eta>0$, there exists $c(\eta)\in(0,\infty)$ so that the following holds. For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a two-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a two-dimensional simple random walk bridge $(S_{t},0\leq t\leq 2n)$ such that
$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|n^{-1/2}S_{2ns}-B_{s}\right|\geq c(\eta)n^{-1/2}\log n\right]\leq c_{0}n^{-\eta}.$
We will be interested in continuous-time random walk loops $\gamma=(S_{t},0\leq t\leq t_{\gamma})$ on $\mathbb{Z}^{d}$. Let $e_{1},\cdots,e_{d}$ be the $d$ unit vectors in $\mathbb{R}^{d}$. Then $S_{t}=S_{t/d}^{1}\,e_{1}+\cdots S^{d}_{t/d}\,e_{d}$ for all $0\leq t\leq t_{\gamma}$, where $S^{1}_{t},\cdots,S^{d}_{t}$ are independent continuous-time random walk bridges on $\mathbb{Z}$. Following the steps of [4], we can similarly deduce the following coupling (we will omit the proof since it is really analogous to that of Theorem 2.2).
Theorem 2.4.
There exist $0<c,\alpha<\infty$, such that for all $t>0$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a continuous-time simple random walk bridge $S$ on $\mathbb{Z}$, both with length $t$, such that for all $r>0$,
$\mathbb{P}\left[\sup_{0\leq s\leq t}|B_{s}-S_{s}|>rc\log t\right]\leq ct^{\alpha-r}.$
This then directly implies the following lemma.
Lemma 2.5.
Fix $d\geq 1$. There exists a constant $c_{0}(d)>0$, such that for each $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds. For every $t>0$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a $d$-dimensional continuous-time simple random walk bridge $(S_{s},0\leq s\leq t)$ such that
$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|(t/d)^{-1/2}S_{st}-B_{s}\right|\geq c(\eta,d)t^{-1/2}\log t\right]\leq c_{0}(d)t^{-\eta}.$
3 Proof of the main theorem
We now construct the coupling and prove our main theorem. We mostly follow the framework of [4], but make some subtle modifications, following a key observation in Section 3.2.
We first treat the general case of continuous-time random walk loop soups on $\mathbb{Z}^{d},d\geq 1$, and then explain in Section 3.4 how to adapt this proof to the special case of discrete random walk loop soups on $\mathbb{Z}$ or $\mathbb{Z}^{2}$.
3.1 Preparation
In this subsection, we define some auxiliary random processes and loops, and construct a family of the random walk loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$, all on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$. We will only look at continuous-time random walk loops, and will omit the word “continuous-time”.
We will construct the random walk loops with time length at least $2$, and then add independently an increasing family of random walk loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ of loops with time length less than $2$. For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, let $N(n,z;t)$ be independent Poisson processes (parametrized by $t$) with parameter $Q_{d}(n)$ given by (2.5). For each $n\geq 1$, $m\geq 1,z\in\mathbb{Z}^{d}$ we define the independent random variables $\widetilde{T}(n,z;m)$ with density
$q_{d}(t)\mathbf{1}_{2n\leq t<2n+2}dt.$
Let $\widetilde{\ell}(n,z;m)$ be a random walk bridge (loop) on $\mathbb{Z}^{d}$ with time length $\widetilde{T}(n,z;m)$. We make the loops $\widetilde{\ell}(n,z;m)$ for different $n,z;m$ independent from each other. Then the collection of loops
$\widetilde{\mathcal{A}}_{\lambda}:=\widetilde{\mathcal{L}}_{\lambda}\cup\big\{\widetilde{\ell}(n,z,m)+z:n\geq 1,z\in\mathbb{Z}^{d},1\leq m\leq N(n,z;\lambda)\big\}$
is a random walk loop soup with intensity $\lambda$. It is increasing in $\lambda$ by construction.
Due to Lemma 2.5, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d. Brownian loops $\ell(n,z;m),n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$, distributed as $d$-dimensional Brownian bridges started at 0 with time length $1$, so that the following holds. By Lemma 2.5, there exist constants $c_{0}(\eta,d)$ and $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-\big(\widetilde{T}(n,z;m)/d\big)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.1)
3.2 Construction of the Brownian loop soup
We now construct a Brownian loop soup on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, using the random processes and loops defined in the previous subsection. The main difference with the proof in [4] is that we introduce an increasing sequence $\{a_{n}\}_{n\geq 1}$, defined by the following recurrence relation
$a_{1}:=(2\pi)^{-1}\bigg((d/2)\int_{2}^{\infty}q_{d}(t)dt\bigg)^{-2/d},\quad a_{n}^{-d/2}-a_{n+1}^{-d/2}=(2\pi)^{d/2}(d/2)Q_{d}(n).$ (3.2)
Note that $q_{d}(t)$ is integrable at $\infty$ by (1.6), so $a_{1}>0$ and the sequence is well defined. This definition ensures that (recall that $(2\pi)^{-d/2}s^{-d/2-1}ds$ is the density of the time length of the Brownian loops in (2.1))
$Q_{d}(n)=\int_{a_{n}}^{a_{n+1}}\frac{ds}{(2\pi)^{d/2}s^{d/2+1}},$ (3.3)
hence we can couple the random walk loops with length in $(n,n+1)$ rooted at $z$, and the Brownian loops with length in $(a_{n},a_{n+1})$ rooted in a unit square centered at $z$, in an exact one-to-one correspondance (i.e. there are the same number of loops in each set).
In order to make our coupling work, we need to first show the following estimate.
Lemma 3.1.
For the sequence $\{a_{n}\}_{n\geq 1}$ defined in (3.2), we have
$a_{n}=2n/d+O(1).$
Proof.
By (3.2), for all $n\geq 2$,
$a_{n}^{-d/2}=a_{1}^{-d/2}-\sum_{k=1}^{n-1}(2\pi)^{d/2}(d/2)Q_{d}(n),$
which tends to 0 as $n\to\infty$, by our choice of $a_{1}$. This implies $a_{n}\to\infty$ as $n\to\infty$.
We also have for $m\geq n\geq 1$,
$a_{n}^{-d/2}-a_{m}^{-d/2}=\sum_{k=n}^{m}(2\pi)^{d/2}(d/2)Q_{d}(n)=(2\pi)^{d/2}(d/2)\int_{2n}^{2m+2}q_{d}(t)dt.$ (3.4)
Letting $m\to\infty$, the left hand-side of (3.4) tends to $a_{n}^{-d/2}$, and the right hand-side of (3.4) tends to $d^{d/2}(2n)^{-d/2}(1+O(n^{-1}))$ by (1.6). This leads to
$a_{n}^{-d/2}=d^{d/2}(2n)^{-d/2}(1+O(n^{-1})).$
It then implies $a_{n}=(2n/d)(1+O(n^{-1}))$, which completes the proof. ∎
We now construct the Brownian loop soup on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$ where the random walk loop soup in Section 3.1 lives. We will only construct the loops with time length at least $a_{1}$, and then add independently an increasing family of Brownian loop soups $\{\mathcal{L}_{\lambda}\}_{\lambda>0}$ of loops with time length less than $a_{1}$. These microscopic loops will not be coupled with the random walk loops.
Recall the Poisson processes $N(n,z;t)$ and the coupled loops $\widetilde{\ell}(n,z;m),\ell(n,z;m)$ defined in Section 3.1. We further define independent complex random variables $Y(n,z;m)$ uniformely distributed on the hypercube $\big\{\sum_{j=1}^{d}x_{j}e_{j},|x_{j}|\leq 1/2\big\}$, and independent real random variables $T(n,z;m)$ with density
$(2\pi)^{-d/2}Q_{d}(n)^{-1}s^{-d/2-1}\mathbf{1}_{a_{n}\leq s\leq a_{n+1}}.$ (3.5)
Indeed, (3.5) is the density of a probability measure by (3.3). We construct the rooted Brownian loop soup as follows
  • Let $N(n,z;t)$ be the number of rooted loops that have appeared by time $t$ whose root is in the unit hypercube centered at $z$ and whose time duration is between $a_{n}$ and $a_{n+1}$;
  • Let $\ell^{*}(n,z;m)$ be the bridge obtained by scaling and translating $\ell(n,z,m)$, so that it has time duration $T(n,z;m)$ and root $z+Y(n,z;m)$.
Then the collection of loops
$\mathcal{A}_{\lambda}:=\mathcal{L}_{\lambda}\cup\{\ell^{*}(n,z,m):n\geq 1,z\in\mathbb{Z}^{d},1\leq m\leq N(n,z;\lambda)\}$
is a Brownian loop soup with intensity $\lambda$. It is increasing in $\lambda$ by construction.
3.3 Proof of Theorem 1.2 for continuous-time random walk loop soups
We will show that the loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ and $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ constructed in the previous subsections satisfy the conditions of Theorem 1.2. Compared with Theorem 1.1, we will make one more (small) improvement. We obtain $c\lambda r^{d}$ as coefficient in the error probability which tends to 0 as $\lambda\to 0$ (compared with the coefficient $c(\lambda+1)r^{2}$ in Theorem 1.1).
Proof of Theorem 1.2.
We fix $a>0$ and $\theta\in(0,2)$. The constants $c_{0},c_{1},\ldots$ that we will use in this proof are allowed to depend on $a,\theta,d$, but not on $r,\lambda,N$. Fix
$k=2+2a/d.$ (3.6)
For $r\geq 1$, $\lambda>0$ and integer $N\geq 1$, let
$Z:=\sum_{|z|\leq rN}\sum_{n\geq N^{k}}N(n,z;\lambda).$
Then $Z$ is a Poisson random variable with
$\mathbb{E}[Z]=\lambda\sum_{|z|\leq rN}\sum_{n\geq N^{k}}Q_{d}(n)=\lambda\sum_{|z|\leq rN}\int_{2N^{k}}^{\infty}q_{d}(t)dt\leq c_{1}\lambda r^{d}N^{d}N^{-kd/2},$
where $c_{1}$ is a constant depending possibly on $a,d$, following (1.6). This implies
$\mathbb{P}\left[N(n,z;\lambda)>0\text{ for some }n\geq N^{k},|z|\leq rN\right]\leq c_{1}\lambda r^{d}N^{(1-k/2)d}=c_{1}\lambda r^{2}N^{-a},$ (3.7)
where the last equality is due to (3.6).
Let us denote the loops $\ell^{*}(n,z;m)$ and $\widetilde{\ell}(n,z;m)+z$ by $\gamma_{n,z,m}$ and $\widetilde{\gamma}_{n,z,m}$. Recall that they have respectively time length $T(n,z;m)$ and $\widetilde{T}(n,z;m)$, which we denote by $T,\widetilde{T}$ for brevity. Fix
$\eta=(a+d)/\theta.$ (3.8)
Recall the constant $c(\eta,d)$ in (3.12). We define the following event
$A:=\bigg\{\sup_{0\leq s\leq 1}$ $|\gamma_{n,z,m}(sT)-\widetilde{\gamma}_{n,z,m}(s\widetilde{T})|\geq 3k\,c(\eta,d)\log N$
$\text{ for some }|z|<rN,N^{\theta}<n<N^{k},1\leq m\leq N(n,z;\lambda)\bigg\}.$
Since
[…]
$+\big|\big((\widetilde{T}/d)^{1/2}T^{-1/2}-1\big)\big(\gamma_{n,z,m}(sT)-Y(n,z;m)\big)\big|$
$+|z-Y(n,z;m)|,$
noting that $|z-Y(n,z;m)|\leq 1$, we have
$A\subseteq\left\{A^{1}_{n,z,m}\cup A^{2}_{n,z,m}\text{ for some }|z|<rN,N^{\theta}<n<N^{k},1\leq m\leq N(n,z;\lambda)\right\},$
where
$A^{1}_{n,z,m}=\bigg\{\sup_{0\leq s\leq 1}\big|(\widetilde{\gamma}_{n,z,m}(s\widetilde{T})-z)-(\widetilde{T}/d)^{1/2}T^{-1/2}(\gamma_{n,z,m}(sT)-Y(n,z;m))\big|\geq c(\eta,d)\log N^{k}\bigg\},$
$A^{2}_{n,z,m}=\bigg\{\big((\widetilde{T}/d)^{1/2}T^{-1/2}-1\big)\sup_{0\leq s\leq 1}|\gamma_{n,z,m}(sT)-Y(n,z;m))|\geq c(\eta,d)\log N^{k}\bigg\}.$
Note that the rescaled loop $T^{-1/2}(\gamma_{n,z,m}(sT)-Y(n,z;m)),0\leq s\leq 1$ is by definition $\ell(n,z;m)$, which is a Brownian bridge of time length $1$. By (3.12), for all $N^{\theta}<n<N^{k}$, there exists a constant $c_{0}$ depending only on $a,\theta,d$ (recall that $\eta$ is a function of $a,\theta$, according to (3.8)), such that
$\mathbb{P}(A^{1}_{n,z,m})\leq c_{0}n^{-\eta}\leq c_{0}N^{-\theta\eta}=c_{0}N^{-a-d}.$ (3.9)
On the other hand, since $2n\leq\widetilde{T}<2n+2$, $a_{n}\leq T\leq a_{n+1}$, and $a_{n}=2n/d+O(1)$ by Lemma 3.1, we have
$\big|(\widetilde{T}/d)^{1/2}T^{-1/2}-1\big|=O(n^{-1}).$
Therefore, there exist $c_{3},c_{4},c_{5}>0$, such that for all $N^{\theta}<n<N^{k}$,
$\mathbb{P}(A^{2}_{n,z,m})\leq\mathbb{P}\bigg[\sup_{0\leq s\leq 1}|\gamma_{n,z,m}(sT)-Y(n,z;m)|\geq c_{3}n\log n\bigg]\leq\exp(-c_{4}n)\leq c_{5}N^{-a-d},$ (3.10)
where the second inequality follows from the fact that $\gamma_{n,z,m}(sT)-Y(n,z;m)$ is a $d$-dimensional Bronwnian bridge with time length $T\sim 2n$, and that the maximum $M$ of a one-dimensional Brownian bridge with time length $1$ satisfies $\mathbb{P}(M>x)=\exp(-2x^{2})$.
Now, let
$W:=\sum_{|z|\leq rN}\sum_{N^{\theta}<n<N^{k}}N(n,z;\lambda).$
Then $W$ is a Poisson random variable with
$\mathbb{E}[W]\leq\lambda\sum_{|z|\leq rN}\sum_{n\geq 1}Q_{d}(n)=c_{6}(d)\lambda r^{d}N^{d},$
where $c_{6}(d)=\sum_{n\geq 1}Q_{d}(n)<\infty$ depends only on $d$. Let $\mathcal{F}_{0}$ be the $\sigma$-algebra generated by $N(n,z;t)$ for all $n\geq 1,z\in\mathbb{Z}^{d},t>0$. Note that $W$ is measurable with respect to $\mathcal{F}_{0}$.
Combining (3.9) and (3.10), we get
$\mathbb{P}(A\mid\mathcal{F}_{0})\leq W(c_{0}+c_{5})N^{-a-d},$
so there exists $c_{7}>0$ such that
$\mathbb{P}(A)\leq\mathbb{E}(W)(c_{0}+c_{5})N^{-a-d}\leq c_{7}\lambda r^{d}N^{-a}.$ (3.11)
On the event $A^{c}\cup\{N(n,z;\lambda)=0\text{ for all }n\geq N^{k},|z|\leq rN\}$ (whose probability is bounded by the sum of (3.7) and (3.11)), the coupling satisfies the conditions of Theorem 1.2, with $c=\max\{c_{1}+c_{7},\,3kc(\eta,d)\}$. Indeed, (1.9) is ensured by $A$, and (1.8) is due to (1.7). ∎
3.4 Proof of Theorem 1.2 for discrete-time random walk loop soups
The proof for discrete-time random walk loop soups on $\mathbb{Z}^{d}$ with $d=1,2$ is similar, so we will only outline the differences.
First, we do not need to add the loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ as in Section 3.1, since there are by definition no loop with time length less than $2$. For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, we similarly define independent Poisson processes $N(n,z;t)$, but with parameter $\widetilde{q}_{d}(n)$ given by (2.2) and (1.5). For each $n\geq 1$, $m\geq 1,z\in\mathbb{Z}^{d}$, let $\widetilde{\ell}(n,z;m)$ be an independent random walk bridge (loop) on $\mathbb{Z}^{d}$ with time length $2n$. Then the collection of loops
$\widetilde{\mathcal{A}}_{\lambda}:=\big\{\widetilde{\ell}(n,z,m)+z:n\geq 1,z\in\mathbb{Z}^{d},1\leq m\leq N(n,z;\lambda)\big\}$
is a random walk loop soup with intensity $\lambda$. It is increasing in $\lambda$ by construction.
Due to Theorem 2.2 and Lemma 2.3, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d. Brownian loops $\ell(n,z;m),n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$, distributed as $d$-dimensional Brownian bridges started at 0 with time length $1$, so that the following holds. By Lemma 2.5, there exist constants $c_{0}(\eta,d)$ and $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-(2n/d)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.12)
We define the sequence $\{a_{n}\}_{n\geq 1}$ by the following recurrence relation
$a_{1}:=(2\pi)^{-1}\bigg((d/2)\sum_{n=1}^{\infty}\widetilde{q}_{d}(n)\bigg)^{-2/d},\quad a_{n}^{-d/2}-a_{n+1}^{-d/2}=(2\pi)^{d/2}(d/2)\widetilde{q}_{d}(n).$ (3.13)
We can then similarly prove that $a_{n}=2n/d+O(1).$ To construct the Brownian loop soup, we proceed exactly as in Section 3.2, but replace $Q_{d}(n)$ by $\widetilde{q}_{d}(n)$ in (3.5).
Acknowledgments
We thank Yifan Gao and Pierre Nolin for useful discussions. We also acknowledge the support by a GRF grant from the Research Grants Council of the Hong Kong SAR (project 11305825).

### v2

@@
 roughly diameter less than $N^{-2/3}$).
-In this paper, we find a simple way to remove the restriction $\theta>2/3$, so that such a coupling works for all $\theta\in(0,2)$, i.e.
+This coupling was subsequently extended in [10] to $\mathbb{Z}^{d}$ with $d\geq 3$, for loops with time length at least $N^{\theta-2}$, for $\theta\in(2d/(d+4),2)$.
+In this paper, we find a simple way to remove the restriction $\theta>2d/(d+4)$, so that such a coupling works for all $\theta\in(0,2)$, i.e.
 for loops at all polynomial scales.
-We also establish this coupling in any dimension $d\geq 1$ (i.e.
-for random walk loop soups on $\mathbb{Z}^{d}$ and Brownian loop soups on $\mathbb{R}^{d}$).
+We establish couplings for both discrete-time and continuous-time random walk loop soups on $\mathbb{Z}^{d}$, for $d\geq 1$.
 1 Introduction
@@
 We then set
-$\mathcal{A}_{\lambda,N}=\{\Phi_{N}\gamma:\gamma\in\mathcal{A}_{\lambda}\}.$
+$\mathcal{A}_{\lambda,N}=\{\Phi_{N}\gamma:\gamma\in\mathcal{A}_{\lambda}\}.$ (1.2)
 The scaling invariance of the Brownian loop soup implies that $\mathcal{A}_{\lambda,N}$ is also distributed as a Brownian loop soup with intensity $\lambda$.
@@
 Theorem 1.1 (Theorem 1.1, [6]).
-One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$, $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{2}$, and $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the random walk loop soup on $\mathbb{Z}^{2}$.
-Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
+One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
 Moreover, for $\mathcal{A}_{\lambda,N}$ and $\widetilde{\mathcal{A}}_{\lambda,N}$ defined as above, there exists $c>0$ such that for every $r\geq 1$, $\lambda>0$, integer $N\geq 1$ and every $\theta\in(2/3,2)$, except perhaps on an event of probability at most $c(\lambda+1)r^{2}N^{2-3\theta}$, there is a one-to-one correspondence between the following two sets of loops
@@
 $|t_{\gamma}-t_{\widetilde{\gamma}}|\leq 5/8N^{-2},\qquad$
-$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$
+$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$ (1.6)
 The coupling by Lawler and Trujillo Ferreras has proved to be useful in a great number of situations, allowing one to transfer the information about the Brownian loop soup back to the random walk loop soup, and vice versa.
@@
 1.2 Main result
-It is unclear at first sight whether $\theta>2/3$ can be removed, [^1] For example, this was believed to be imposssible in [1, Section 11.3]: “On the other hand, it is fairly clear from the proof of [4] that their result is sharp, and that the coupling described above cannot hold without the restriction $\theta>2/3$; that is, at all scales smaller than $N^{-2/3}$ some discrete and continuous loops somewhere will be quite different from one another.” since the value $2/3$ comes from a computation involving the following Taylor expansion of the total mass $q_{2}(n)$ of all random walk loops on $\mathbb{Z}^{2}$ with length $2n$ rooted at 0
+It is unclear at first sight whether $\theta>2/3$ can be removed, [^1] For example, this was believed to be impossible in [1, Section 11.3]: “On the other hand, it is fairly clear from the proof of [6] that their result is sharp, and that the coupling described above cannot hold without the restriction $\theta>2/3$; that is, at all scales smaller than $N^{-2/3}$ some discrete and continuous loops somewhere will be quite different from one another.” since the value $2/3$ comes from a computation involving the following Taylor expansion of the total mass $q_{2}(n)$ of all random walk loops on $\mathbb{Z}^{2}$ with length $2n$ rooted at 0
 $\widetilde{q}_{2}(n)=(2n)^{-1}\left[2^{-2n}\binom{2n}{n}\right]^{2}=\frac{1}{2\pi n^{2}}-\frac{1}{8\pi n^{3}}+O(n^{-4}).$ (1.7)
@@
 Our observation is rather simple, but it seems that it has been overlooked.
-We will directly establish our result for random walks on $\mathbb{Z}^{d}$ for all $d\geq 1$.
-Note that the combinatorial formula (1.5) relies on a trick which is specific for $d=2$, namely we can rotate $\mathbb{Z}^{2}$ by $45$ degrees, so that the original simple random walk projected to the vertical and horizontal axes are independent simple random walks.
-For $d\geq 3$, it is more convenient to consider the continuous-time simple random walk, which spends an (independent) exponential time with mean $1$ at each vertex.
-We will rely on the following Taylor development of the total mas $q_{d}(t)$ of all continuous-time random walk loops on $\mathbb{Z}^{d}$ with time length $t$ rooted at 0 (see (2.3) and (2.4))
+Lawler and Trujillo Ferreras established this coupling in dimension $2$, but their proof has been extended to dimensions $d\geq 3$ by Sapozhnikov and Shiraishi in [10, Theorem 2.2], relying on the expansion (see e.g.
+[2])
+$\widetilde{q}_{d}(n)=2\bigg(\frac{d}{4\pi n}\bigg)^{d/2}\bigg(1-\frac{d}{8n}+O\big(\frac{1}{n^{2}}\big)\bigg),$ (1.8)
+where $\widetilde{q}_{d}(n)$ is the total mass of all random walk loops on $\mathbb{Z}^{d}$ with length $2n$ rooted at 0.
+Similarly, the coupling in [10] only works for loops with length greater than $N^{\theta-2}$, where
+$2d/(d+4)<\theta<2,$ (1.9)
+on an event with probability at least
+$1-c(\lambda+1)r^{d}N^{-\min\left(\frac{d}{2},\;\theta\left(\frac{d}{2}+2\right)-d\right)}$ (1.10)
+for some $c\in(0,\infty)$.
+They also obtained a slightly weaker bound $cN^{-1/4}\log N$ on the supreme distance between the coupled loops, instead of $cN^{-1}\log N$ in (1.6).
+This is because they established a weaker coupling between random walk bridges on $\mathbb{Z}^{d}$ with the Brownian bridges (see Lemma 2.3), for $d\geq 3$.
+Note that the strong KMT coupling in dimension $2$ established in [6] (see Lemma 2.2), as well as the exact combinatorial formula (1.7), relies on a trick which is specific to $d=2$: We can rotate $\mathbb{Z}^{2}$ by $45$ degrees, so that the original simple random walk projected to the vertical and horizontal axes are independent simple random walks.
+In this paper, we consider both the discrete-time random walk loop soups (as in [6, 10]), and the continuous-time random walk loop soups on $\mathbb{Z}^{d}$, where every walk spends an (independent) exponential time with mean $1$ at each vertex.
+The latter version is more natural when one considers the occupation time field of the loop soup, due to a well-known isomorphism theorem [8].
+In addition, the latter version is equal to the “restriction” of the Brownian loop soup on the cable graph of $\mathbb{Z}^{d}$ (i.e.
+we restrict the Brownian loops on the cable graph to stay on a vertex, before jumping to the next vertex).
+The two versions of random walk loop soups are the same, if one forgets the time parametrization and remove the single-vertex loops in the continuous-time random walk loop soup (see [5, Section 9]).
+However, since our coupling takes into account the time-parametrization of the loops, the couplings for the two versions of loop soups need separate proofs, which can be done in parallel (also see Remark 1.5).
+By abuse of notation, we use $\widetilde{\mathcal{A}}_{\lambda}$ to denote both versions of random walk loop soups.
+The actual version we use will be specified in the context.
+The continuous-time random walk has the advantage that its movements in the $d$ coordinates are independent, hence we can obtain the same strong KMT coupling as in [6], yielding the same $cN^{-1}\log N$ bound as in (1.6) on the supreme distance between the coupled loops.
+In this case, we will rely on the following Taylor development of the total mas $q_{d}(t)$ of all continuous-time random walk loops on $\mathbb{Z}^{d}$ with time length $t$ rooted at 0 (see (2.3) and (2.4))
 $q_{d}(t)=(d/2)^{d/2}\pi^{-d/2}t^{-d/2-1}+O\big(t^{-d/2-2}\big).$ (1.11)
-Our proof requires only the first term in the Taylor series, unlike (1.5).
+Our proof requires only the first term in the Taylor series, unlike in [6, 10].
 In order to state our result, we keep the same notations as in Section 1.1, with the following obvious extensions.
@@
 For $t\geq a_{1}N^{-2}$ and integer $k\geq 1$, let
-$\chi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{a_{k}}{N^{2}}\leq t<\frac{a_{k+1}}{N^{2}}.$
+$\chi_{N}(t)=\frac{k}{N^{2}}\quad\text{if}\quad\frac{a_{k}}{N^{2}}\leq t<\frac{a_{k+1}}{N^{2}}.$ (1.13)
 Theorem 1.2.
-Fix $d\geq 1$. One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that for each $\lambda>0$,
-  1. (1)
-    $\mathcal{A}_{\lambda}$ is a realization of the Brownian loop soup in $\mathbb{R}^{d}$;
-  2. (2)
-    $\widetilde{\mathcal{A}}_{\lambda}$ is a realization of the continuous-time random walk loop soup on $\mathbb{Z}^{d}$ (or $\widetilde{\mathcal{A}}_{\lambda}$ can also be a discrete-time random walk loop soup if $d=1,2$).
-Both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
+Fix $d\geq 1$.
+Let $\mathcal{A}_{\lambda}$ be a Brownian loop soup on $\mathbb{R}^{d}$, and $\widetilde{\mathcal{A}}_{\lambda}$ be a (continuous-time) random walk loop soup on $\mathbb{Z}^{d}$.
+One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
 Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
@@
 $\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N.$ (1.15)
-This strengthened coupling, which controls loops at any polynomial scale, rooted anywhere in a macroscopic domain, can have many potential applications.
+Note that Theorem 1.2 works for loops at any polynomial scale, rooted anywhere in a macroscopic domain.
+Moreover, the error probability (of the event on which the coupling fails) is at most $c\lambda r^{d}N^{-a}$, where $a$ can be made arbitrarily large, compared with Theorem 1.1 and (1.10).
+This strengthened coupling can have many potential applications.
 For example, in the two-dimensional case, it was proved by [9, 12] that the outer boundaries of the outermost clusters of the random walk loop soups with intensity $\lambda\in(0,1/2]$ converge to those of the Brownian loop soups (which are in turn distributed as conformal loop ensembles [11]).
 This convergence is nevertheless qualitative.
-If one wants to obtain a precise convergence rate, one may need to control the random walk loops at mesoscopic scales, everywhere in the domain simultaneously, since the outer boundary of a cluster is a global object.
-On the other hand, for dimensions $d\geq 3$, to our knowledge, the convergence of random walk loop soups in $\mathbb{Z}^{d}$ to Brownian loop soups has not previously been written down, so we fill in this small gap in the literature.
-Note that it is clear that the same coupling also works for Brownian loop soups on the metric graph induced by $\mathbb{Z}^{d}$ (since its restriction to the vertices of $\mathbb{Z}^{d}$ is exactly a continuous-time random walk loop soup).
-This type of convergence has been used in e.g.
-[10, 7].
+If one wants to obtain a precise convergence rate, it then seems necessary to control the random walk loops at mesoscopic scales, everywhere in the domain simultaneously, since the outer boundary of a cluster is a global object.
+The result for discrete-time random walk loop soups on $\mathbb{Z}^{d}$ is rather similar, but we still state it below in Theorem 1.3, so that the reader can directly compare it with [6] in the $d=2$ case (and with [10] in the $d\geq 3$ case).
+In the discrete-time case, we use a different sequence $\{a_{n}\}_{n\geq 1}$ defined in (3.14), which also satisfies (1.12).
+We keep the same definition (1.13) of $\chi_{N}$, but inputting the new sequence $\{a_{n}\}_{n\geq 1}$.
+Theorem 1.3.
+Fix $d\geq 1$.
+Let $\mathcal{A}_{\lambda}$ be a Brownian loop soup on $\mathbb{R}^{d}$, and $\widetilde{\mathcal{A}}_{\lambda}$ be a (discrete-time) random walk loop soup on $\mathbb{Z}^{d}$.
+One can define $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ and $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ on the same probability space so that both $\mathcal{A}_{\lambda}$ and $\widetilde{\mathcal{A}}_{\lambda}$ are increasing in $\lambda$.
+Moreover, for every $a>0$ and every $\theta\in(0,2)$, there exists $c>0$ such that for every $r\geq 1,\lambda>0$ and integer $N\geq 1$, except perhaps on an event of probability at most $c\lambda r^{d}N^{-a}$, there is a one-to-one correspondence between
+$\{\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}:t_{\widetilde{\gamma}}>N^{\theta-2},|\widetilde{\gamma}(0)|<r\},\quad\{\gamma\in\mathcal{A}_{\lambda,N}:\chi_{N}(t_{\gamma})>N^{\theta-2},|\psi_{N}(\gamma(0))|<r\}.$
+If $\widetilde{\gamma}\in\widetilde{\mathcal{A}}_{\lambda,N}$ and $\gamma\in\mathcal{A}_{\lambda,N}$ are paired in this correspondence, then for the constant $c(d)$ in (1.12),
+$|t_{\gamma}-t_{\widetilde{\gamma}}|\leq c(d)N^{-2},\qquad$ (1.16)
+$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{-1}\log N\qquad\text{ if }d=1,2,$ (1.17)
+$\sup_{0\leq s\leq 1}$ $|\gamma(st_{\gamma})-\widetilde{\gamma}(st_{\widetilde{\gamma}})|\leq cN^{\frac{a-d}{2d}}\log N\quad\;\text{ if }d\geq 3.$ (1.18)
+Remark 1.4.
+For $d\geq 3$, we get the bound $cN^{\frac{a-d}{2d}}\log N$ in (1.18), using the coupling between the random walk bridge and the Brownian bridge (Lemma 2.3) established in [10].
+This bound tends to 0 as $N\to\infty$ only when $a<d$, hence we cannot reduce the error probability (of the event on which the coupling fails) below $c\lambda r^{d}N^{-d}$ in this case.
+Note that this does not prevent us from choosing $\theta$ close to 0, hence this coupling does still control loops at all polynomial scales.
+If we choose $a=d/2$, then we get the same $cN^{-1/4}\log N$ bound as in [10].
+We believe that it is possible to improve the bound (1.18), by improving the coupling of Lemma 2.3.
+However, if we do not care about the time parametrization, for example if we only care about the Hausdorff distance between the coupled loops, then (1.15) already provides the strong bound $cN^{-1}\log N$.
+Remark 1.5.
+An immediate corollary of Theorems 1.2 and 1.3 is that the discrete-time and continuous-time random walk loop soups can also be coupled together, since they can both be coupled with the Brownian loop soup.
+This coupling is, however, quite different from the obvious coupling where we let the two versions of loop soups have the same trace on $\mathbb{Z}^{d}$ (except for the single-vertex loops).
+Note that two random walk loops with the same trace, respectively parametrized by discrete and continuous time, typically do not have time lengths close to each other in the sense of (1.16).
 2 Preliminaries
+In this section, we recall the definitions and some basic properties of the Brownian loop soup, the discrete-time and continuous-time random walk loop soups on $\mathbb{Z}^{d}$, and record several versions of couplings between Brownian bridges and random walk bridges.
 2.1 Brownian loop soup
@@
 Let (note that $|\nu_{n,z}|$ is the same for all $z$)
-$\widetilde{q}_{d}(n):=|\nu_{n,z}|.$
-For $d=1,2$, we have the exact combinatorial formula (1.5) for $\widetilde{q}_{2}(n)$, and
-$\widetilde{q}_{1}(n)=(2n)^{-1}2^{-2n}\binom{2n}{n}$. (2.2)
+$\widetilde{q}_{d}(n):=|\nu_{n,z}|.$ (2.2)
+The combinatorial formula of $\widetilde{q}_{2}(n)$ is given in (1.7) (and a similar formula can easily be obtained for $\widetilde{q}_{1}(n)$).
+In general, the Taylor development of $\widetilde{q}_{d}(n)$ is given by (1.8).
 2.3 Continuous-time random walk loop soup
@@
 $q_{d}(t):=\frac{1}{t}p_{t}(0,0).$ (2.3)
-Thanks to the following Taylor development of $p_{t}(0,0)$ (see e.g. [3, Theorem 2.13])
+Thanks to the following Taylor development of $p_{t}(0,0)$ (see e.g. [5, Theorem 2.13])
 $p_{t}(0,0)=\bigg(\frac{d}{2\pi t}\bigg)^{d/2}(1+O(t^{-1})),$ (2.4)
@@
 $Q_{d}(n):=\int_{2n}^{2n+2}q_{d}(t)dt.$ (2.5)
-2.4 KMT coupling
-To couple a Brownian loop closely to a random walk loop, we rely on a variant of the classical coupling by Komlós, Major, and Tusnády [2].
-Recall that the KMT coupling works for a wide class of random walks, and a comprehensive review of it can be found in [3, Chapter 7].
-In particular, for any aperiodic irreducible walk $S_{n}=X_{1}+\cdots+X_{n}$ on $\mathbb{Z}$ such that there exist $\sigma,b>0$ with
-$\mathbb{E}[X_{1}]=0,\quad\mathbb{E}[X_{1}^{2}]=\sigma^{2},\quad\mathbb{E}[\exp(b|X_{1}|)]<\infty,$ (2.6)
-we can couple it with a Brownian motion (see e.g. [3, Theorem 7.1.1]).
+2.4 Coupling Brownian bridges and random walk bridges
+In order to couple a Brownian bridge closely to a random walk bridge on the same probability space, we rely on several variants of the classical coupling by Komlós, Major, and Tusnády [4].
+We will consider discrete-time random walk bridges in Theorem 2.1, Lemmas 2.2 and 2.3, and continuous-time random walk bridges in Theorem 2.4 and Lemma 2.5.
+For both cases, we regard the random walk bridges as continuous curves, by linear interpolation.
+First of all, we need the following coupling between a discrete-time simple random walk bridge on $\mathbb{Z}$ and a Brownian bridge with length $2n$, established in [6, Theorem 6.4].
 Theorem 2.1.
-Suppose that $(S_{n})_{n\geq 0}$ is an aperiodic irreducible random walk on $\mathbb{Z}$ that satisfies (2.6).
-Then one can define on the same probability space a Brownian motion $B_{t}$ with variance $\sigma^{2}$ and $(S_{n})_{n\geq 0}$ such that the following holds.
-For each $\alpha<\infty$, there is a $c_{\alpha}$ such that
-$\mathbb{P}\left[\max_{1\leq j\leq n}|S_{j}-B_{j}|\geq c_{\alpha}\log n\right]\leq c_{\alpha}n^{-\alpha}$
-The idea of the proof of Theorem 2.1 has been used in [4] to prove the following coupling between a discrete-time simple random walk bridge on $\mathbb{Z}$ a Brownian bridge with length $2n$ (see [4, Theorem 6.4]).
-Note that even though the random walk bridges in Theorem 2.2 and Lemma 2.3 below are defined on discrete time and space, we still consider them as continuous curves parametrized by continuous time intervals, by linear interpolation.
-This is, however, different from continuous-time random walks.
-Theorem 2.2.
-There exist $0<c,\alpha<\infty$, such that for every positive integer $n$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a simple random walk bridge $S$ on $\mathbb{Z}$, both with length $2n$, such that for all $r>0$,
+There exist $0<c,\alpha<\infty$, such that for every positive integer $n$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a (discrete-time) random walk bridge $S$ on $\mathbb{Z}$, both with length $2n$, such that for all $r>0$,
 $\mathbb{P}\left[\sup_{0\leq t\leq 2n}|B_{t}-S_{t}|>rc\log n\right]\leq cn^{\alpha-r}.$
-By the trick of rotating $\mathbb{Z}^{2}$ by $45$ degrees (see the second paragraph of Section 1.2), the simple random walk on $\mathbb{Z}^{2}$ is the combination of two independent simple random walks on $\mathbb{Z}$, hence Theorem 2.2 implies the following coupling (see [4, Lemma 3.1] and the remark just below it).
+By the trick of rotating $\mathbb{Z}^{2}$ by $45$ degrees (see the second paragraph of Section 1.2), the simple random walk on $\mathbb{Z}^{2}$ is the combination of two independent simple random walks on $\mathbb{Z}$, hence Theorem 2.1 implies the following coupling (see [6, Lemma 3.1] and the remark just below it).
 Lemma 2.2.
 There exists a constant $c_{0}>0$, such that for each $\eta>0$, there exists $c(\eta)\in(0,\infty)$ so that the following holds.
-For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a two-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a two-dimensional simple random walk bridge $(S_{t},0\leq t\leq 2n)$ such that
-$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|n^{-1/2}S_{2ns}-B_{s}\right|\geq c(\eta)n^{-1/2}\log n\right]\leq c_{0}n^{-\eta}.$
-We will be interested in continuous-time random walk loops $\gamma=(S_{t},0\leq t\leq t_{\gamma})$ on $\mathbb{Z}^{d}$.
+For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a two-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a two-dimensional (discrete-time) random walk bridge $(S_{t},0\leq t\leq 2n)$ such that
+$\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|n^{-1/2}S_{2ns}-B_{s}\right|\geq c(\eta)n^{-1/2}\log n\right]\leq c_{0}n^{-\eta}.$ (2.6)
+In [10, Lemma 4.2], the authors established the following coupling between random walk bridges on $\mathbb{Z}^{d}$ and $d$-dimensional Brownian bridges, for $d\geq 3$.
+They got a weaker $n^{-1/4}$ factor in (2.7), compared with $n^{-1/2}$ in (2.6), due to the technical difficulty arising from the fact that the movements of the random walk in $d$ coordinates are not independent.
+Lemma 2.3.
+Fix $d\geq 3$.
+For any $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds.
+For every positive integer $n$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a (discrete-time) random walk bridge $(S_{t},0\leq t\leq 2n)$ on $\mathbb{Z}^{d}$ such that
+$\mathbb{P}\left[\sup_{0\leq s\leq 1}\big|(2n/d)^{-1/2}S_{2ns}-B_{s}\big|>c(\eta,d)n^{-1/4}\log n\right]\leq c(\eta,d)n^{-\eta}.$ (2.7)
+We now turn to the continuous-time random walk bridge on $\mathbb{Z}$.
+Its coupling with the Brownian bridge is analogous to the discrete-time case, and has been proved in [3], where the authors actually considered random bridges constructed through a rather general class of random walks with i.i.d.
+jumps.
+Let us state the following version that we will use.
+Theorem 2.4.
+There exist $0<c,\alpha<\infty$, such that for all $t>0$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a (continuous-time) random walk bridge $S$ on $\mathbb{Z}$, both with length $t$, such that for all $r>0$,
+$\mathbb{P}\left[\sup_{0\leq s\leq t}|B_{s}-S_{s}|>rc\log t\right]\leq ct^{\alpha-r}.$
+We now consider a continuous-time random walk bridge $\gamma=(S_{t},0\leq t\leq t_{\gamma})$ on $\mathbb{Z}^{d}$.
 Let $e_{1},\cdots,e_{d}$ be the $d$ unit vectors in $\mathbb{R}^{d}$.
 Then $S_{t}=S_{t/d}^{1}\,e_{1}+\cdots S^{d}_{t/d}\,e_{d}$ for all $0\leq t\leq t_{\gamma}$, where $S^{1}_{t},\cdots,S^{d}_{t}$ are independent continuous-time random walk bridges on $\mathbb{Z}$.
-Following the steps of [4], we can similarly deduce the following coupling (we will omit the proof since it is really analogous to that of Theorem 2.2).
-Theorem 2.4.
-There exist $0<c,\alpha<\infty$, such that for all $t>0$, there is a probability space on which are defined a one-dimensional Brownian bridge $B$ and a continuous-time simple random walk bridge $S$ on $\mathbb{Z}$, both with length $t$, such that for all $r>0$,
-$\mathbb{P}\left[\sup_{0\leq s\leq t}|B_{s}-S_{s}|>rc\log t\right]\leq ct^{\alpha-r}.$
-This then directly implies the following lemma.
+Theorem 2.4 then directly implies the following lemma.
 Lemma 2.5.
@@
 There exists a constant $c_{0}(d)>0$, such that for each $\eta>0$, there exists $c(\eta,d)\in(0,\infty)$ so that the following holds.
-For every $t>0$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a $d$-dimensional continuous-time simple random walk bridge $(S_{s},0\leq s\leq t)$ such that
+For every $t>0$, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which are defined a $d$-dimensional Brownian bridge $(B_{t},0\leq t\leq 1)$ and a $d$-dimensional (continuous-time) random walk bridge $(S_{s},0\leq s\leq t)$ such that
 $\mathbb{P}\left[\sup_{0\leq s\leq 1}\left|(t/d)^{-1/2}S_{st}-B_{s}\right|\geq c(\eta,d)t^{-1/2}\log t\right]\leq c_{0}(d)t^{-\eta}.$
-3 Proof of the main theorem
+3 Proof of the main theorems
 We now construct the coupling and prove our main theorem. We mostly follow the framework of [6], but make some subtle modifications, following a key observation in Section 3.2.
-We first treat the general case of continuous-time random walk loop soups on $\mathbb{Z}^{d},d\geq 1$, and then explain in Section 3.4 how to adapt this proof to the special case of discrete random walk loop soups on $\mathbb{Z}$ or $\mathbb{Z}^{2}$.
+We first treat, in Sections 3.1, 3.2 and 3.3, the case of continuous-time random walk loop soups.
+We omit the adjective “continuous-time” in our description of random walk loop soups in these sections.
+We then explain, in Section 3.4, how to adapt this proof to the case of discrete-time random walk loop soups.
 3.1 Preparation
 In this subsection, we define some auxiliary random processes and loops, and construct a family of the random walk loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$, all on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$.
-We will only look at continuous-time random walk loops, and will omit the word “continuous-time”.
 We will construct the random walk loops with time length at least $2$, and then add independently an increasing family of random walk loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ of loops with time length less than $2$.
@@
 By Lemma 2.5, there exist constants $c_{0}(\eta,d)$ and $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
-$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-\big(\widetilde{T}(n,z;m)/d\big)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.1)
+$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(sT)-\big(\widetilde{T}/d\big)^{1/2}\ell(n,z;m)(s)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta},$ (3.1)
+where we write $T=T(n,z;m)$ and $\widetilde{T}=\widetilde{T}(n,z;m)$ for brevity.
 3.2 Construction of the Brownian loop soup
@@
 Recall the Poisson processes $N(n,z;t)$ and the coupled loops $\widetilde{\ell}(n,z;m),\ell(n,z;m)$ defined in Section 3.1.
-We further define independent complex random variables $Y(n,z;m)$ uniformely distributed on the hypercube $\big\{\sum_{j=1}^{d}x_{j}e_{j},|x_{j}|\leq 1/2\big\}$, and independent real random variables $T(n,z;m)$ with density
+We further define independent complex random variables $Y(n,z;m)$ uniformly distributed on the hypercube $\big\{\sum_{j=1}^{d}x_{j}e_{j},|x_{j}|\leq 1/2\big\}$, and independent real random variables $T(n,z;m)$ with density
 $(2\pi)^{-d/2}Q_{d}(n)^{-1}s^{-d/2-1}\mathbf{1}_{a_{n}\leq s\leq a_{n+1}}.$ (3.5)
@@
 is a Brownian loop soup with intensity $\lambda$. It is increasing in $\lambda$ by construction.
-3.3 Proof of Theorem 1.2 for continuous-time random walk loop soups
+3.3 Proof of Theorem 1.2
 We will show that the loop soups $\{\widetilde{\mathcal{A}}_{\lambda}\}_{\lambda>0}$ and $\{\mathcal{A}_{\lambda}\}_{\lambda>0}$ constructed in the previous subsections satisfy the conditions of Theorem 1.2.
@@
 where $c_{1}$ is a constant depending possibly on $a,d$, following (1.11). This implies
-$\mathbb{P}\left[N(n,z;\lambda)>0\text{ for some }n\geq N^{k},|z|\leq rN\right]\leq c_{1}\lambda r^{d}N^{(1-k/2)d}=c_{1}\lambda r^{2}N^{-a},$ (3.7)
+$\mathbb{P}\left[N(n,z;\lambda)>0\text{ for some }n\geq N^{k},|z|\leq rN\right]\leq c_{1}\lambda r^{d}N^{(1-k/2)d}=c_{1}\lambda r^{d}N^{-a},$ (3.7)
 where the last equality is due to (3.6).
@@
 $\mathbb{P}(A^{2}_{n,z,m})\leq\mathbb{P}\bigg[\sup_{0\leq s\leq 1}|\gamma_{n,z,m}(sT)-Y(n,z;m)|\geq c_{3}n\log n\bigg]\leq\exp(-c_{4}n)\leq c_{5}N^{-a-d},$ (3.10)
-where the second inequality follows from the fact that $\gamma_{n,z,m}(sT)-Y(n,z;m)$ is a $d$-dimensional Bronwnian bridge with time length $T\sim 2n$, and that the maximum $M$ of a one-dimensional Brownian bridge with time length $1$ satisfies $\mathbb{P}(M>x)=\exp(-2x^{2})$.
+where the second inequality follows from the fact that $\gamma_{n,z,m}(sT)-Y(n,z;m)$ is a $d$-dimensional Brownian bridge with time length $T\sim 2n$, and that the maximum $M$ of a one-dimensional Brownian bridge with time length $1$ satisfies $\mathbb{P}(M>x)=\exp(-2x^{2})$.
 Now, let
@@
 ∎
-3.4 Proof of Theorem 1.2 for discrete-time random walk loop soups
-The proof for discrete-time random walk loop soups on $\mathbb{Z}^{d}$ with $d=1,2$ is similar, so we will only outline the differences.
-First, we do not need to add the loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ as in Section 3.1, since there are by definition no loop with time length less than $2$.
-For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, we similarly define independent Poisson processes $N(n,z;t)$, but with parameter $\widetilde{q}_{d}(n)$ given by (2.2) and (1.5).
+3.4 Proof of Theorem 1.3
+The proof for discrete-time random walk loop soups is similar, so we will only outline the differences.
+For simplicity, we keep the same notations as for the continuous-time random walk loop soups, but update some of their definitions.
+First, we do not need to add the loop soups $\{\widetilde{\mathcal{L}}_{\lambda}\}_{\lambda>0}$ as in Section 3.1, since there are by definition no loop with time length less than $2$ in the discrete-time random walk loop soup.
+For integers $n\geq 1$ and $z\in\mathbb{Z}^{d}$, we similarly define independent Poisson processes $N(n,z;t)$, but with parameter $\widetilde{q}_{d}(n)$ given by (2.2), instead of $Q_{d}(n)$.
 For each $n\geq 1$, $m\geq 1,z\in\mathbb{Z}^{d}$, let $\widetilde{\ell}(n,z;m)$ be an independent random walk bridge (loop) on $\mathbb{Z}^{d}$ with time length $2n$.
@@
 is a random walk loop soup with intensity $\lambda$. It is increasing in $\lambda$ by construction.
-Due to Theorem 2.2 and Lemma 2.3, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d.
+For $d=1,2$, due to Theorem 2.1 and Lemma 2.2, we can also construct on the same probability space $(\Omega,\mathcal{F},\mathbb{P})$, i.i.d.
 Brownian loops $\ell(n,z;m),n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$, distributed as $d$-dimensional Brownian bridges started at 0 with time length $1$, so that the following holds.
-By Lemma 2.5, there exist constants $c_{0}(\eta,d)$ and $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
-$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)-(2n/d)^{1/2}\ell(n,z,m)\big|\geq c(\eta,d)\log n\bigg]\leq c_{0}(\eta,d)n^{-\eta}.$ (3.12)
+$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(2ns)-(2n/d)^{1/2}\ell(n,z,m)(s)\big|\geq c(\eta,d)\log n\bigg]\leq c(\eta,d)n^{-\eta}.$ (3.12)
+For $d\geq 3$, by Lemma 2.3, we get a slightly weaker bound, namely there exist $c(\eta,d)$, such that for all $n\geq 1,z\in\mathbb{Z}^{d},m\geq 1$,
+$\mathbb{P}\bigg[\sup_{0\leq s\leq 1}\big|\widetilde{\ell}(n,z;m)(2ns)-(2n/d)^{1/2}\ell(n,z,m)(s)\big|\geq c(\eta,d)n^{1/4}\log n\bigg]\leq c(\eta,d)n^{-\eta}.$ (3.13)
+The bounds (3.12) and (3.13) are the analogues of (3.1).
 We define the sequence $\{a_{n}\}_{n\geq 1}$ by the following recurrence relation
@@
 We can then similarly prove that $a_{n}=2n/d+O(1).$ To construct the Brownian loop soup, we proceed exactly as in Section 3.2, but replace $Q_{d}(n)$ by $\widetilde{q}_{d}(n)$ in (3.5).
+For $d=1,2$, the rest of the proof is exactly the same as in Section 3.3. For $d\geq 3$, we still need to update the definitions of the events $A$, $A^{1}_{n,z,m}$ and $A^{2}_{n,z,m}$ to
+$\begin{split}A:=\bigg\{\sup_{0\leq s\leq 1}&|\gamma_{n,z,m}(sT)-\widetilde{\gamma}_{n,z,m}(s\widetilde{T})|\geq 3k\,c(\eta,d)N^{k/4}\log N\\ &\text{ for some }|z|<rN,N^{\theta}<n<N^{k},1\leq m\leq N(n,z;\lambda)\bigg\}\end{split}$ (3.15)
+and
+$A^{1}_{n,z,m}=\bigg\{\sup_{0\leq s\leq 1}\big|(\widetilde{\gamma}_{n,z,m}(s\widetilde{T})-z)-(\widetilde{T}/d)^{1/2}T^{-1/2}(\gamma_{n,z,m}(sT)-Y(n,z;m))\big|\geq c(\eta,d)N^{k/4}\log N^{k}\bigg\},$
+$A^{2}_{n,z,m}=\bigg\{\big((\widetilde{T}/d)^{1/2}T^{-1/2}-1\big)\sup_{0\leq s\leq 1}|\gamma_{n,z,m}(sT)-Y(n,z;m))|\geq c(\eta,d)N^{k/4}\log N^{k}\bigg\}.$
+By (3.13), we deduce for all $N^{\theta}<n<N^{k}$, there exists a constant $c_{0}$ depending only on $a,\theta,d$, such that $\mathbb{P}(A^{1}_{n,z,m})\leq c_{0}N^{-a-d}$.
+We get a similar bound for $\mathbb{P}(A^{2}_{n,z,m})$, as in (3.10).
+The definition (3.15) of $A$ implies the bound (1.18) for the loops which are rescaled by $N^{-1}$ in space, noting that $k=2+2a/d$ by (3.6).
 Acknowledgments
-We thank Yifan Gao and Pierre Nolin for useful discussions. We also acknowledge the support by a GRF grant from the Research Grants Council of the Hong Kong SAR (project 11305825).
+We thank Yifan Gao, Xinyi Li and Pierre Nolin for useful discussions. We also acknowledge the support by a GRF grant from the Research Grants Council of the Hong Kong SAR (project 11305825).